What mass of sodium oxide,Na2O will be obtain by burning 11.5g of sodium in air
1
Expert's answer
2021-09-27T01:48:39-0400

4Na+O2=2Na2O

v(Na)=11.5g/23g/mol=0.5mol

v(Na2O)=v/2(Na)=0.25mol

m(Na2O)=0.25mol*62g/mol=15.5g